OpenStudy (anonymous):

Determine how much money will be in a savings account if the initial deposit was $1000 and the interest rate is 2.467% compounded continuously for 25 years. (Round your answer to the nearest cent.)

OpenStudy (anonymous):

$1000e^(2.467(25))

OpenStudy (anonymous):

For some reason when I type it into my calculator, it doesn't work

OpenStudy (lukebluefive):

That's not quite right. It's actually: \[1000 \times e^\left( 0.02467 \times 25 \right)\] Your calculator is most likely choking because that other number is really large.

OpenStudy (anonymous):

Oh, sorry I forgot to convert it to a decimal. Thank you for the catch, Luke.

OpenStudy (lukebluefive):

@ash.vasvani np, you did most of the work. Thanks for the assist!

OpenStudy (anonymous):

Thank you both of you guys!